# Generalizable Articulated Object Reconstruction from Casually Captured RGBD Videos

### v1

@@ 4 Experiments
We aim to answer the following questions with our experiments: 1) How well can we estimate joint parameters? 2) How well can we reconstruct the geometry of the articulated object? 3) How do different modules and strategies of our pipeline contribute to the final results? and 4) Can our method work in real-world scenarios?
@@ 4.1 Experiment setup
At present, there is no benchmark for evaluating joint parameter estimation from dynamic RGBD videos. Therefore, we construct our synthetic dataset for evaluation. We choose objects from 11 categories in the PartNet-Mobility dataset [25]. We load them into the SAPIEN simulator [49] to simulate interaction between humans and objects. For each interaction, we generate two videos shot from two different viewpoints. To generate camera motion, every 7 - 10 timesteps in the simulator, the camera will move to a random target pose. The target pose is sampled from a Gaussian distribution, where the mean value is the current camera pose and the standard deviation is set to ensure the camera can consistently look at the object. For the object surface point cloud, we simply fuse 24 views’ point clouds of the object. For Robot-See-Robot-Do, we generate 100 views because it requires to train NeRF and 3D gaussian splatting [10]. Please refer to Appendix B for more details about our dataset. Since Robot-See-Robot-Do requires manually cropping and clustering 3D gaussian splatting [10], which is labor-intensive, we split the dataset into two parts. We randomly sampled 10% of test videos from the whole dataset to construct a smaller dataset named S-Dataset, containing 73 test videos, while L-Dataset contains the remaining videos in the original dataset.
Baselines
We choose two very recent works on modeling and reconstructing articulated objects that take videos as input:
@@
Robot-See-Robot-Do (RSRD): The authors introduce a differentiable rendering pipeline to recover 3D part motion from human demonstration videos [10]. We follow their pipeline first to reconstruct and segment the 3D gaussian splatting of the object. Then, input a video to estimate the SE3 transformation of each part of 3D gaussians at each frame via differentiable rendering. Since their pipeline does not include geometry reconstruction module, we export 3D gaussians from their pipeline for geometry reconstruction evaluation for this method.
@@
Articulate Anything: The authors introduce a system to retrieve object mesh from a library and then estimate link and joint parameters from text, image, or video input [13]. We follow their setting to use the Partnet-Mobility dataset as the object mesh library, which contains all the ground truth object meshes for our test cases. For VLM, we also follow their setting to use Gemini Flash-1.5 model [37].
Metrics
For articulated object reconstruction, we evaluate the following metrics: 1) Joint parameters, including joint type classification error in percentage, joint axis error in randian, joint position error for revolute joint in meter, and joint state error in randian for revolute joint and meter for prismatic joint. 2) Chamfer distance on the whole object, movable part of the object, and static part, denoted as CD-w, CD-m, CD-s, respectively. If the method crashes or retrieves the wrong objects, we classify these cases as prediction failures and assign for joint axis and state of revolute joint error, and 1 for the other metrics. Due to the page limit, we put more results and discussions in Appendix C
4.2 Joint estimation results
Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad/m) | Failure(%) |
Revolute | Ours | 0.320.56 | 0.130.25 | 15.90 | 0.250.46 | 6.81 |
Articulate Anything | 0.820.79 | 0.810.40 | 40.90 | N.A. | 38.63 | |
RSRD | 1.170.51 | 2.037.48 | 88.63 | 1.030.60 | 47.72 | |
Ours w/o Refine | 1.030.67 | 0.240.25 | 68.18 | 0.590.47 | 4.54 | |
Ours w/o Coarse | 1.260.29 | 0.340.26 | 77.27 | 0.690.45 | 2.27 | |
Ours w/o Segment | 0.360.55 | 0.170.24 | 18.18 | 0.330.44 | 6.81 | |
Prismatic | Ours | 0.240.33 | N.A. | 10.34 | 0.080.22 | 0 |
Articulate Anything | 0.920.78 | N.A. | 50.00 | N.A. | 46.66 | |
RSRD | 1.220.42 | N.A. | 41.37 | 0.700.48 | 37.93 | |
Ours w/o Refine | 0.280.40 | N.A. | 6.89 | 0.160.16 | 0 | |
Ours w/o Coarse | 0.420.11 | N.A. | 6.89 | 0.080.09 | 0 | |
Ours w/o Segment | 0.300.41 | N.A. | 13.79 | 0.070.07 | 0 |
@@ 4.2 Joint estimation results
RSRD | 3.3921.47 | 0.821.17 | 0.140.41 |
Ours w/o Refine | 0.010.01 | 0.400.44 | 0.350.46 |
Ours w/o Coarse | 0.010.01 | 0.190.22 | 0.060.16 |
Ours w/o Segment | 0.010.01 | 0.400.44 | 0.350.46 |
@@
We first comprehensively evaluate our method and two baselines on the S-Dataset. The quantitative results on S-Dataset are in Tab. 1, and qualitative results are in Fig. 4. Results on L-Dataset can be found in Sec. C.1. Our proposed method is comparable and superior in most metrics to the current state-of-the-art method on our dataset. In our dataset, Articulate Anything struggles with retrieving the correct object meshes. Most failure cases in Tab. 1 are due to retrieval error, even though the library contains all the ground truth object meshes. For example, in Fig. 4 we find that Articulate Anything retrieves a lamp in the Stapler test case. Besides, the link placement and affordance detection modules are not robust either. In Fig. 4, Articulate Anything selects the wheel of the table to be the movable part in the Table test case, and places the lid of the USB below the main body in the USB test case. For RSRD, we spot 3D gaussian splatting in their pipeline cannot represent objects with less texture information accurately. Therefore, it often fails to crop and cluster 3D gaussians correctly, though human operators participate in this step. Besides, their pipeline cannot handle camera motion. Thus, its prediction deviates from the ground truth a lot as shown in Fig. 4.
4.3 Geometric reconstruction results
@@
We evaluate the geometric reconstruction results with the Chamfer distance. Since mesh is the target geometry representation of the articulated objects, we sample 10000 points from the reconstructed mesh and the ground truth mesh, and compute the Chamfer distance between them. Tab. 2 shows the results of geometry reconstruction on S-Dataset. Results on L-Dataset can be found in Sec. C.1. Due to the retrieval error, link placement error, and affordance detection error, Articulate Anything has a high reconstruction error, particularly on movable parts. RSRD performs even worse on geometry reconstruction since it struggles with clustering 3D gaussian splatting into different parts.
4.4 Ablation study
@@
We design three ablations of our method. To demonstrate the importance of the coarse prediction module, we remove the coarse prediction module from our pipeline and randomly initialize the parameters of the refinement module, named Ours w/o Coarse. To illustrate the importance of the refinement module, we remove the refinement module and evaluate coarse prediction results directly, named Ours w/o Refine. Lastly, to demonstrate the advantage of the automatic part segmentation strategy for the refinement process, we naively optimize the probability of being a moving part for each pixel at each video frame, named Ours w/o Segment.
@@
The joint parameter prediction results are shown in Tab. 1. Without refinement, coarse prediction will fail on joint type prediction for more than half of the test cases, particularly on revolute joints. As we mentioned in Sec. 3, when two video frames are similar, it’s hard to distinguish revolute joint and prismatic joint. In Fig. 4, we can also find that for Stapler and Microwave test cases, the coarse prediction module misclassifies these two test cases. Besides, without the refinement module, it’s much harder to correctly segment the object’s movable parts and static parts, as shown in Tab. 2. Without coarse prediction, refinement cannot optimize effectively with random initialization, as shown quantitatively in Tab. 2 and Tab. 1 and qualitatively in Fig. 4. Without automatic part segmentation, the joint prediction performance degrades as the refinement module suffers from the incorrect point correspondence issue mentioned in Sec. 3. For instance, it fails on the Microwave test case in Fig. 4. Besides, the segmentation also deviates from the correct segmentation, as shown in Tab. 2. In Fig. 4, we can also find that without automatic part segmentation, our method has difficulty segmenting the moving parts and static parts accurately. This illustrates the importance of automatic part segmentation in optimizing joint parameters and moving map of the video efficiently.
@@ 4.5 Real-world qualitative evaluation
We also evaluate our methods and baselines on real data. We use an iPhone 12 Pro with LiDAR camera to capture real data. We first use Polycam [28] to reconstruct the object’s surface, and then use Record3D [31] to record an RGB-D video of human interaction with the object. To further improve the depth map quality, we adopt Prompt Depth Anything [15] to scale up the original depth map. Real videos are likely to include human hands, which are other moving objects in the scene. This will cause confusion to MonST3R when predicting the moving map. Thus, we use Grounded SAM 2 [32, 33] to mask hands with simple text prompts: "hands and arms".
From the qualitative results shown in Fig. 5, Articulate Anything fails on the Book test case due to the lack of similar meshes in the mesh library. In the Cabinet test case, Articulate Anything retrieves reasonable meshes but predicts the wrong joint rotation direction in the Cabinet test case, as the door rolls inside the cabinet. RSRD does not work in either case. This demonstrates our method can work on real data and outperform current methods. We put more qualitative results in Sec. C.7.
@@ 5 Conclusion
We introduced a new problem setting for reconstructing articulated objects from casually captured RGBD videos. Our problem setting is more practical than prior works, bringing new challenges to current methods. We then developed a coarse-to-fine framework for this problem. To evaluate, we proposed a new dataset that contains 20 more objects than previous benchmarks. Our experiments show our method outperforms baselines, providing a practical and robust approach to reconstructing articulated objects. Nonetheless, there is still space for improvement. Better dynamic scene understanding approaches and video segmentation models can help improve our method. In addition, our pipeline does not reconstruct the interior of the object. Incrementally reconstructing the interior of the object is another future direction.
Acknowledgments and Disclosure of Funding
@@ A.1 Method
Since the input video usually contains more than 100 frames, we sample the input video to around 20 frames. We first send these sampled frames to MonST3R [55] to compute the moving map of each frame. Then, we select pairs of frames to compute the feature matching using LoFTr [35]. Not all feature matches are accurate, so we only consider feature matches that have a confidence value larger than 0.95 to be valid. We first use feature matching within the static regions to compute the relative camera poses from frame to sequentially. We use these relative camera poses to align all the frames to the first camera coordinate. Then, we proceed to estimate joint parameters. We select frame pairs within three intervals to balance the accuracy of feature matching and joint prediction. Due to the inaccuracy of moving map and feature matching, we filter out frame pairs that contain fewer than 80 valid feature matches. For the rest of the frame pairs, we estimate joint parameters for the revolute joint and the prismatic joint separately. For each joint type, we run RANSAC [4] for 50 iterations to select the joint parameters estimation that best fits the observation. For each frame pair, we also compare the distances between corresponding feature points under two sets of joint parameters. If the distance for the revolute joint is lower, this frame pair votes for revolute, and vice versa. After estimating joint parameters for all the frame pairs, we compute the mean value for the joint axis and position for both joint types. We choose the joint type that gets the most votes as the predicted joint type for coarse prediction.
Refinement
@@
Though we predict the joint type in coarse prediction, we still retain the prediction results of both joint types from coarse prediction and initialize the refinement module with those values. For moving vector initialization, we compute the intersection area between each part and the moving map from MonST3R [55], and then normalize it by the area of each part. During the optimization, we compute the optimization objective according to Eqs. 1 and 2. But computing the chamfer distance will be very slow if the number of points is very high. Therefore, for each frame, we randomly sample 50% of points from the whole observation to compute the Chamfer distance. The gradient will not be affected since we are optimizing the moving vector instead of the probability of being a moving part of each pixel at each frame. This strategy significantly improves the optimization speed without losing prediction accuracy. We concurrently optimize the joint prediction for both joint types for 400 iterations, using Adam optimizer [11] and 5e-3 learning rate. After optimization, we need to choose the joint type. We first classify the test case as a prismatic joint if the distance between the object surface point cloud and the joint axis is larger than 0.1 meters. For the remaining cases, we compare the lowest chamfer distance of these two cases and select the joint type that has the lower chamfer distance. These are the final joint type prediction results for our method. Finally, we find pixels in the moving map whose value is larger than 0.7 to be the moving points. We project these points back to 3D and find points in the object surface point cloud whose distance to the moving points is less than 0.03 meters to be the movable parts of the object. The other points in the object surface point cloud are classified as static parts of the object. We reconstruct meshes using NKSR [6].
@@ Appendix B Dataset details
We build our dataset in the SAPIEN simulator [49]. We design a small cuboid space with a simple texture as the environment. Then, we place the object in that space and render videos while manipulating the object using the built-in controller. Fig. B.1 illustrates the data collection environment. We build a dataset for robust evaluation of our method in Tab. B.1.
@@ C.1 Joint estimation and geometry reconstruction results on L-Dataset
Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad/m) | Failure(%) |
Revolute | Ours | 0.290.51 | 0.120.22 | 14.80 | 0.270.46 | 3.43 |
Articulate Anything | 0.730.77 | 0.730.40 | 37.12 | N.A. | 35.19 | |
Ours w/o Refine | 1.090.62 | 0.240.33 | 73.60 | 0.590.51 | 2.57 | |
Ours w/o Coarse | 1.300.25 | 0.340.23 | 76.82 | 0.680.44 | 1.07 | |
Ours w/o Segment | 0.490.62 | 0.200.22 | 26.60 | 0.380.45 | 3.21 | |
Prismatic | Ours | 0.270.40 | N.A. | 8.16 | 0.060.15 | 2.44 |
Articulate Anything | 0.830.78 | N.A. | 49.79 | N.A. | 44.89 | |
Ours w/o Refine | 0.270.44 | N.A. | 8.57 | 0.100.08 | 0.40 | |
Ours w/o Coarse | 0.510.20 | N.A. | 14.69 | 0.120.15 | 1.63 | |
Ours w/o Segment | 0.290.42 | N.A. | 9.79 | 0.080.15 | 2.44 |
@@
Ours w/o Refine | 0.010.03 | 0.320.42 | 0.270.42 |
Ours w/o Coarse | 0.010.03 | 0.150.25 | 0.050.14 |
Ours w/o Segment | 0.010.03 | 0.350.43 | 0.300.44 |
We conduct a more extensive evaluation of methods, excluding RSRD, on the L-Dataset. Quantitative results on joint estimation and geometry reconstruction are shown in Tab. C.2 and Tab. C.3. Similar to the results in S-Dataset, our method outperforms Articulate Anything and other ablated versions consistently on L-Dataset.

### v2

@@ 4 Experiments
We aim to answer the following questions with our experiments: 1) How well can we estimate joint parameters? 2) How well can we acquire the geometry of the articulated object? 3) How do different modules and strategies of our pipeline contribute to the final results? and 4) Can our method work in real-world scenarios?
@@ 4.1 Experiment Setup
At present, there is no benchmark for evaluating joint parameter estimation from dynamic RGBD videos. Therefore, we construct a synthetic dataset for evaluation. We select objects from 11 categories in the PartNet-Mobility dataset [33]. We simulate interaction between human hands and these objects using the SAPIEN simulator [67]. For each interaction, we generate two videos shot from two different viewpoints. Camera trajectories are generated by moving the camera towards a new target pose every 7 - 10 timesteps in the simulator. The target pose is sampled from a Gaussian distribution, where the mean value is the current camera pose and the standard deviation is set to ensure the camera can consistently look at the object. To get the object surface point cloud, we unproject the depth maps from 24 viewpoints to 3D and fuse the point cloud. For Robot-See-Robot-Do, we generate 100 views because it requires producing a NeRF and 3D Gaussian splat [14]. Please refer to the supplement for more details about the dataset construction. Since Robot-See-Robot-Do requires manually cropping and clustering 3D gaussian splats which is labor-intensive, we split the dataset into two parts. We randomly sampled 10% of test videos from our dataset to construct a smaller dataset named S-Dataset, containing 73 test videos, while L-Dataset contains the remaining videos in the original dataset.
Baselines.
We select two recent works on modeling interactable digital twins of articulated objects from video input:
@@
Robot-See-Robot-Do (RSRD) [14]: a differentiable rendering pipeline that recovers 3D part motion from human demonstration videos. We follow the pipeline to first reconstruct and segment the 3D Gaussian splats of the object. Then, the input video is used to estimate the SE3 transformation of the 3D Gaussians for each part at each frame via differentiable rendering. For a fair comparison, we modify their tracking pipeline to exploit ground truth depth maps of the input video. We select the group of 3D Gaussians with the longest moving distance to be the moving part of the object. Since this pipeline does not include a geometry reconstruction module, we export 3D Gaussians and reconstruct mesh using NKSR [9] for geometry reconstruction.
@@
Articulate Anything [18]: a system that retrieves object meshes from a library and estimates link and joint parameters from text, image, or video input. We follow the original work and use the Partnet-Mobility dataset as the object mesh library, which contains all the ground truth object meshes for our test cases, and use Gemini Flash-1.5 [51] for the VLM module.
Metrics.
We use a suite of metrics to evaluate both geometric reconstruction and articulation parameter estimation. Geometric accuracy is evaluated using Chamfer Distance between ground truth and estimated geometry for the whole object (CD-w), for the movable part (CD-m) and for the static part (CD-s). Articulation parameter estimation is evaluated using joint type classification error (percentage), joint axis error (radians), joint position error for revolute joints (meters), and joint state error (radians for revolute joints and meters for prismatic joints). If the method crashes or retrieves incorrect objects (for methods based on retrieval), we classify the output as a failure and assign error for the joint axis and state of revolute joints, and error equal to 1 for the other metrics.
4.2 Articulation Parameter Estimation Results
@@
| RSRD | 3.3921.50 | 0.600.60 | 0.170.44 |
| iTACO (Ours) | 0.010.01 | 0.130.26 | 0.060.19 |
| iTACO w/o Refine | 0.010.01 | 0.400.44 | 0.350.46 |
| iTACO w/o Coarse | 0.010.01 | 0.190.22 | 0.060.16 |
| iTACO w/o Segment | 0.010.01 | 0.390.44 | 0.360.46 |
@@
The quantitative results on S-Dataset are in Tab.˜1, and qualitative results are in Fig.˜4. Results on L-Dataset can be found in the supplement. Our iTACO method outpeforms or is comparable to state-of-the-art methods along most metrics. Articulate Anything struggles with retrieving the correct object meshes. Most failure cases in Tab.˜1 are due to retrieval errors, which is surprising since the retrieval library contains all the ground truth object meshes. For example, in Fig.˜4 we find that Articulate Anything retrieves a lamp in the Stapler test case. In addition, the link placement and affordance detection modules appear to not be robust either, showing the challenge of our problem setting. Articulate Anything selects the wheel of the table to be a movable part, and places the lid of the USB stick below the main body.
At the same time, RSRD cannot represent textureless objects accurately, likely due to limitations in using 3D Gaussian splatting. Therefore, it often fails to correctly crop and cluster 3D Gaussians into parts, even though human operators participate in this step. Most importantly, RSRD cannot handle camera motion. Thus, its movable part segmentation deviates from the ground truth significantly. In contrast, iTACO produces more accurate object articulations.
4.3 Geometry Estimation Results
@@
We evaluate the geometry of our outputs using the Chamfer Distance to the ground truth. We sample 10000 points from the reconstructed mesh and the ground truth mesh, and compute the Chamfer Distance between them. Tab.˜2 shows the results onon S-Dataset. Results on L-Dataset are in the supplement. Due to retrieval errors, link placement errors, and affordance detection errors, Articulate Anything has a high reconstruction error, particularly on movable parts. RSRD performs even worse on since it struggles with clustering 3D Gaussian splats into different parts.
4.4 Ablations
@@
We design three ablations of our method. To demonstrate the importance of the coarse prediction module, we remove it from our pipeline and randomly initialize the parameters of the refinement module, named iTACO w/o Coarse. To illustrate the importance of the refinement module, we remove the refinement module and evaluate coarse prediction results directly, named iTACO w/o Refine. Lastly, to demonstrate the advantage of the automatic part segmentation strategy for the refinement process, we naïvely optimize the probability of being a moving part for each pixel at each video frame, named iTACO w/o Segment.
@@
In Tab.˜1 we find that without refinement, coarse prediction will fail on joint type prediction for more than half of the test cases, particularly on revolute joints. As mentioned in Sec.˜3, when two video frames are similar it is hard to distinguish revolute joints from prismatic joints. In Fig.˜4, we also see that the coarse prediction module misclassifies the Stapler and Microwave joints. Without the refinement module it is much harder to correctly segment the movable and static parts, as shown in Tab.˜2. Also, refinement cannot work effectively from a random initialization in the ablation without coarse prediction, as shown quantitatively in Tabs.˜2 and 1 and qualitatively in Fig.˜4.
Without automatic part segmentation, the joint prediction performance degrades as the refinement module suffers from incorrect point correspondences. For instance, the Microwave test case in Fig.˜4 fails. The segmentation also deviates from the correct segmentation as shown in Tab.˜2. In Fig.˜4, we also see that without automatic part segmentation, our method has difficulty segmenting the moving parts and static parts accurately. This illustrates the importance of automatic part segmentation in optimizing joint parameters and the moving map from the video.
@@ 4.5 Real-World Qualitative Evaluation
We also evaluate our methods and baselines on real data. We use an iPhone 12 Pro with LiDAR camera to capture data. We use Polycam [37] to reconstruct the object surface and Record3D [40] to record the RGBD video of the human interaction with the object. To improve depth map quality, we use Prompt Depth Anything [22] to scale up the original depth map. We compute the image feature matching between the first frame of the video and all the images of the reconstruction stage with LoFTr [46]. We select the image pair with the most reliable feature matching, and compute the rigid transformation between corresponding points of the image pair to align the first video frame to the object surface coordinate system. Real videos include human hands, which interfere with MonST3R predictions for the moving map. Thus, we use Grounded SAM 2 [41, 42] to mask hands with the text prompt: “hands and arms”.
From the qualitative results shown in Fig.˜5, we see that Articulate Anything fails on the Book test case due to the absence of similar meshes in the mesh library. In the Cabinet test case, Articulate Anything retrieves reasonable meshes but predicts the wrong joint rotation direction, as the door rotates into the cabinet. RSRD fails to produce valid results due to the clustering and tracking issues on 3D gaussian. This demonstrates our method can work on real data and outperform current methods in the challenging setting of casually captured real-world RGBD video inputs. See the supplement for more results.
@@ 5 Conclusion
We introduced a new problem setting for reconstructing articulated objects from casually captured RGBD videos. Our problem setting is more practical than prior works, bringing new challenges to current methods. We then developed iTACO: a coarse-to-fine framework for this problem. To evaluate our method, we collected a new dataset that contains 20 more objects than previous benchmarks. Our experiments show our method outperforms baselines, providing a practical and robust approach to acquiring digital twins of articulated objects. Nonetheless, there is still space for improvement. Better dynamic scene understanding approaches and video segmentation models can help improve our method. In addition, our pipeline does not reconstruct the interior of the object. Incrementally reconstructing the interior of the object is an interesting direction for future work.
Acknowledgement
@@ A.1 Method
Since the input video usually contains more than 100 frames, we sample the input video to around 20 frames. We first send these sampled frames to MonST3R [82] to compute the moving map of each frame. Then, we select pairs of frames to compute the feature matching using LoFTr [46]. Not all feature matches are accurate, so we only consider feature matches that have a confidence value larger than 0.95 to be valid. We first use feature matching within the static regions to compute the relative camera poses from frame to sequentially. We use these relative camera poses to align all the frames to the first camera coordinate.
Then, we proceed to estimate joint parameters. We select frame pairs within 3 time steps to ensure the accuracy of feature matching. Due to the inaccuracy of moving map and feature matching, we filter out frame pairs that contain fewer than 80 valid feature matches. For the rest of the frame pairs, we estimate joint parameters for the revolute joint and the prismatic joint separately. For each joint type, we run RANSAC [7] for 50 iterations to select the joint parameters estimation that best fits the observation. Based on the estimated joint parameters, we can transform the feature points to expect position at time step . By computing the distance between the expected position and the observed position , we can evaluate how well this set of joint parameters describes the observation. If the distance for the revolute joint is lower, this frame pair votes for revolute, and vice versa. After estimating joint parameters for all the frame pairs, we compute the mean value for the joint axis and position for both joint types. We choose the joint type that gets the most votes as the predicted joint type for coarse prediction.
Refinement.
@@
Though we predict the joint type in coarse prediction, we still retain the prediction results of both joint types from coarse prediction and initialize the refinement module with those values separately. For moving vector initialization, we compute the intersection area between each part and the moving map from MonST3R [82], and then normalize it by the area of each part. We run the forward computation described in the main paper. But computing the Chamfer Distance will be very slow if the number of points is very high. Therefore, for each frame, we randomly sample 50% of points from the whole observation to compute the Chamfer Distance. The gradient will not be affected since we are optimizing the moving vector instead of the probability of being a moving part of each pixel at each frame. This strategy significantly improves the optimization speed without losing prediction accuracy. We concurrently optimize the joint prediction for both joint types for 400 iterations, using Adam optimizer [15] and 5e-3 learning rate.
After optimization, we have two sets of optimized joint parameters, and we need to choose the joint type. We first classify the test case as a prismatic joint if the distance between the object surface point cloud and the joint axis is larger than 0.1 meters. For the remaining cases, we compare the lowest chamfer distance of these two joint types and select the joint type that has the lower Chamfer Distance. These are the final joint type prediction results for our method. Finally, we find pixels in the moving map whose value is larger than 0.7 to be the moving points. We project these points back to 3D and find points in the object surface point cloud whose distance to the moving points is less than 0.03 meters to be the movable parts of the object. The other points in the object surface point cloud are classified as static parts of the object. We reconstruct meshes using NKSR [9].
Hardware Requirements.
@@ Appendix B Dataset Details
We build our dataset in the SAPIEN simulator [67]. We design a small cuboid space with a simple texture as the environment. Then, we place the object in that space and render videos while manipulating the object using the built-in controller. Fig.˜B.1 illustrates the data collection environment. The statistics are listed in Tab.˜B.1.
@@ C.1 Joint Estimation and Geometry Reconstruction Results on L-Dataset
| Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad or m) | Failure(%) |
| Revolute | Articulate Anything | 0.730.77 | 0.730.40 | 37.12 | N.A. | 35.19 |
| iTACO (Ours) | 0.290.51 | 0.120.22 | 14.80 | 0.270.46 | 3.43 | |
| iTACO w/o Refine | 1.090.62 | 0.240.33 | 73.60 | 0.590.51 | 2.57 | |
| iTACO w/o Coarse | 1.300.25 | 0.340.23 | 76.82 | 0.680.44 | 1.07 | |
| iTACO w/o Segment | 0.490.62 | 0.200.22 | 26.60 | 0.380.45 | 3.21 | |
| Prismatic | Articulate Anything | 0.830.78 | N.A. | 49.79 | N.A. | 44.89 |
| iTACO | 0.270.40 | N.A. | 8.16 | 0.060.15 | 2.44 | |
| iTACO w/o Refine | 0.270.44 | N.A. | 8.57 | 0.100.08 | 0.40 | |
| iTACO w/o Coarse | 0.510.20 | N.A. | 14.69 | 0.120.15 | 1.63 | |
| iTACO w/o Segment | 0.290.42 | N.A. | 9.79 | 0.080.15 | 2.44 |
@@
| iTACO (Ours) | 0.010.03 | 0.100.21 | 0.060.19 |
| iTACO w/o Refine | 0.010.03 | 0.320.42 | 0.270.42 |
| iTACO w/o Coarse | 0.010.03 | 0.150.25 | 0.050.14 |
| iTACO w/o Segment | 0.010.03 | 0.350.43 | 0.300.44 |
We conduct a more extensive evaluation of methods, excluding RSRD, on the L-Dataset. Quantitative results on joint estimation and geometry reconstruction are shown in Tab.˜C.2 and Tab.˜C.3. Similar to the results in S-Dataset, our method outperforms Articulate Anything and other ablated versions consistently on L-Dataset.